OpenStudy (anonymous):

A skier moves 100.0 m horizontally, and then travels another 35.0 m uphill at an angle of 35.0º above the horizontal. What is the skier's displacement from his starting point? Give both the magnitude and the angle in your answer.

OpenStudy (anonymous):

i understand that but then i dont understand how to find the angle. i think the hyp side is 42 or 47. but im not sure

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

\[\sin\theta=\frac{\text{opposite side}}{\text{hypotenuse}}\] \[{\text{hypotenuse}}\times\sin\theta={\text{opposite side}}\]

OpenStudy (unklerhaukus):

35 m is the hypotenuse